xx TestCoordinates

TestCoordinates

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 bitantxplane10 Scope: restricted01  KEYWORD

11 Description: Plane defining bitant domain

11 Range 10                01  Default: xy
11 xy  10 xy-plane
11 xz  10 xz-plane
11 yz  10 yz-plane

11

11 domain  10 Scope: restricted01  KEYWORD

11 Description: Domain type

11 Range 10                01  Default: full
11 octant  10 Use an octant about the origin
11 quadrant 10 Use a quadrant in x-y plane
11 bitant  10 Use a bitant about the x-y plane
11 full  10 Use the full domain

11

7 Interfaces

General Implements: testgrid

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  cartesian  10                    01  compact 11 0
11                    10  x  01 description 11 Cartesian coordinates
11                    10  y  01 dimensions 11 3
11                    10  z  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  spherical  10                    01  compact 11 0
11                    10  r  01 description 11 Spherical polar coordinates
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestCoordinates. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 cartesian spherical 11 x

11 x  11 x

11

Scheduled Functions

CCTKxSTARTUP x coordinatesxstartup x register coordinates x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxBASEGRID x coordinatesxsetup x set up coordinates x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxEVOL x coordinatesxquery x query coordinates x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function